### v1

# Orca: Browsing at Scale Through User-Driven and AI-Facilitated Orchestration Across Malleable Webpages

Abstract.
Web-based activities are fundamentally distributed across webpages. However, conventional browsers with stacks of tabs fail to support operating and synthesizing large volumes of information across pages. While recent AI systems enable fully automated web browsing and information synthesis, they often diminish user agency and hinder contextual understanding. Therefore, we explore how AI could instead augment users’ interactions with content across webpages and mitigate cognitive and manual efforts. Through literature on information tasks and web browsing challenges, and an iterative design process, we present a rich set of novel interactions with our prototype web browser, Orca. Leveraging AI, Orca supports user-driven exploration, operation, organization, and synthesis of web content at scale. To enable browsing at scale, webpages are treated as malleable materials that humans and AI can collaboratively manipulate and compose into a malleable, dynamic, and browser-level workspace. Our evaluation revealed an increased “appetite” for information foraging, enhanced user control, and more flexibility in sensemaking across a broader information landscape on the web.
Web browser, Malleable user interface, Web automation, Large language model
ccs: Human-centered computing Human computer interaction (HCI)ccs: Information systems Information systems applicationsccs: Information systems Browsers
Figure 1. The Orca browser enables browsing at scale by supporting viewing (1), navigating (2), organizing (3), extracting (4), operating (5), and synthesizing at scale.
1. Introduction
Despite the growing scale and complexity of the web and the information tasks we perform on it, the web browser—the primary interface between users and the web—has seen little fundamental change over the past two decades. The conventional linear stack of tabbed pages has not only fallen short in easing the burden of discovering and reviewing large volumes of information during foraging but also adds considerable overhead in managing and synthesizing information scattered across multiple tabs (Chang et al., 2021a; Ma et al., 2023).
Recent advances in Large Language Models (LLMs) have brought transformative changes to how we interact with information on the web. For example, OpenAI’s Operator integrates a chat interface with a classic browser, which helps translate users’ natural language prompts into a sequence of operations on a browser (ope, 2025). Systems like Deep Research further remove the user from the browsing and synthesis process, delivering a final report (dee, 2025). While these approaches can reduce the cognitive and manual effort of browsing, managing, and synthesizing content, direct user interaction with web content remains irreplaceable for several reasons. First, compared to AI-generated summaries, directly engaging with content allows users to consume information in its original context, enabling a more accurate understanding of the problem space (Pirolli and Card, 1995). Second, user-driven browsing lets users stay in control of information sources, steer their evolving exploration directions, preserve a sense of agency, and leave room for serendipity (Foster and Ford, 2003; Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005).
Therefore, instead of using AI to take the users away from the web, we aim to explore enhancing user-driven browsing by preserving user engagement with web content and employing AI to mitigate undesired cognitive and manual effort. We explore how browsers can be redesigned to enable browsing at scale. Our key approach is to treat webpages as “malleable materials”—elements that can be viewed in parallel and flexibly extracted—and the browser as a “malleable space”–an environment where these materials can be dynamically reorganized and composed into a custom workspace. With AI facilitation, users can then view, organize, and manipulate content across webpages in a scalable and flexible manner, which we define as “orchestration.”
To ground our design exploration that enables browsing at scale, we synthesized the literature that identified the merits of manual effort and cognitive costs during web browsing and sensemaking. Reflecting on the activities involved—from starting to synthesizing—we break down the broader goal of browsing-at-scale into specific interaction challenges and opportunities.
Building on these insights, we present a rich set of novel interactions resulting from our exploration with Orca  [^1] Resources are available at https://hci.ucsd.edu/orca and https://jiang.pl/orca., a prototype web browser. Orca supports user-driven exploration, operation, organization, and synthesis of web content at scale while leveraging AI to help users stay in control. For example, with Orca, users can open a large number of links in batches to broaden the exploration, with autonomous agents concurrently deployed across tabs, to accelerate repetitive interactions occurring across webpages. Webpages can be transformed into more succinct forms and flexibly organized into different structures based on their content and users’ context. Finally, information across pages can be synthesized on demand, supporting progressive sensemaking for the dynamic webpages and exploration. As a result, Orca helps users build and manage a dynamic information space, with the web as its foundation.
To understand how effectively Orca supports user-driven browsing, we conducted a lab study with eight participants. We found that Orca not only accelerated web exploration, but also encouraged broader and more confident foraging by reducing the overhead of managing context across pages. Participants valued the parallel view for maintaining awareness and organizing tasks, and appreciated the ability to selectively delegate to AI while retaining control over information sources. These findings suggest that Orca facilitates user-driven, at-scale browsing by aligning AI assistance with users’ evolving goals, preferences, and sensemaking processes.
2. Problem Framing
This work emerges amid rapid advancements in LLMs and their applications. Fully automated information agents like OpenAI Deep Research, Perplexity, and Manus now demonstrate considerable capabilities to forage information, synthesize, and generate comprehensive reports from simple user queries, transforming how we access the information (Venkit et al., 2024; dee, 2025; per, 2022; man, 2025).
However, HCI and sensemaking research have long emphasized the benefits of the information foraging and synthesizing process itself (Rieh et al., 2016; Teevan et al., 2004). Therefore, this work takes a different stance—We do not want to replace browsing with fully automated agents. Instead, we explore leveraging AI’s capabilities and integrating them into each activity in the information-seeking and sensemaking process, to reduce cognitive and interaction costs while preserving the cognitive benefits of direct information engagement.
To capture users’ informational needs and both the challenging and rewarding aspects of user-driven browsing, we review literature on information gathering, sensemaking, and web-based information tasks. We structure our problem framing by adapting Ellis’ behavioral model of information retrieval to include synthesis and transactional activities, seeking the full range of modern web tasks (Table 1(Ellis, 1989; Ellis et al., 1993; Ellis and Haugan, 1997; Choo et al., 1999; Chang et al., 2021a; Sellen et al., 2002). For each activity, we identify costs, highlight cognitive benefits of user-driven browsing, and propose opportunities for AI facilitation.
2.1. Starting
Informational tasks often start without well-defined goals and domain-specific knowledge and language that are needed to translate users’ vague needs into concrete queries (Aula and Russell, 2008; Marchionini, 2006; Hearst, 2009; Ellis, 1989). As a result, selecting starting points and crafting initial queries can be demanding and often leads to confusion and ineffective paths, requiring users to iteratively adjust their queries and sources throughout the process (Aula and Russell, 2008; Bates, 1989; White and Roth, 2009; Kuhlthau, 1991). On the other hand, being able to flexibly restart and reformulate queries allows users to adaptively anchor and reorient their exploration and gain a deeper understanding of their goal and necessary vocabulary (Vakkari, 2001; Rieh et al., 2014).
Different approaches for query assistance and recommendation have been explored to help people formulate and reformulate queries to search more effectively, including assisting completion and refinement (Palani et al., 2021; Lau and Horvitz, 1999; Agapie et al., 2013), session-based search (Radlinski et al., 2010), and related query recommendations (Jones et al., 2006; Kelly et al., 2009; Silvestri, 2009; Williams, 1984). Contextual information, such as browsing history and notes have been leveraged to generate relevant suggestions (Palani et al., 2021; Chang et al., 2019). However, prior work focuses mainly on query assistance within scoped search engines or sources. Despite reviewing results helping users clarify their goals and improve queries, they still have to choose from recommendations and perform each search manually and repeatedly.
In this work, we explore moving beyond simple query suggestions by enabling AI to proactively search multiple sources in parallel, helping users quickly access, overview, and compare results across sources to support broader exploration.
Table 1. Activities, Costs, and Benefits of User-Driven Browsing, and Opportunities for Facilitation
Activities Tasks Goals Costs to Mitigate Benefits to Preserve Opportunities
Starting Section 2.1 Formulate query, select source Establish an initial direction for search Repeatedly translate goals to queries and compare results among different sources Enable goal-oriented and anchored exploration Bootstrap explorations through parallel starting points across sources at scale
Scouting Section 2.2 Navigate links and citations, scan search results and lists Expand, reorient, and contextualize the scope of search Handling multiple threads across isolated pages causes information overload and context loss Enable contextual understanding and serendipitous discovery Accelerate exploration and cross-reference through navigation and overview at scale
Filtering Section 2.2 Remove irrelevant pages or content, highlight and bookmark favorites Identify and prioritize relevant items while discarding the rest Navigate through a large amount of irrelevant information and materials Refine understanding and criteria that lead to more accurate future queries Support evaluating, comparing, and refining exploration results at scale
Revisiting Section 2.3 Organize and revisit content, track changes Organize, track, and revisit useful content Locating stacked pages and tracking scattered information are effortful and disorienting Refresh memories of information and form new connections and ideas with updated context Flexibly organize webpages at scale to support refind and diverse task needs
Collecting Section 2.3 Save, copy, or import information to workspaces Capture, transform, and aggregate content across sources Saving information from multiple disconnected sources continuously disrupts workflow Externalize understanding and build structured, curated knowledge Extract and overview information across webpages at scale
Synthesizing Section 2.4 Compare, integrate, and schematize information Integrate and structure materials into a coherent understanding Integrating fragmented content into coherent understanding is cognitively demanding Connect insights across sources and deepen understanding Dynamically synthesize web content across webpages as scale
Transacting Section 2.4 Complete operations and communicating information Act upon, interpret, and communicate the findings Performing repetitive actions across interfaces is tedious and time-consuming Build user agency and trust towards the final results Streamline repetitive operations across pages at scale through parallel, on-demand automation
2.2. Scouting and Filtering
When an information task begins, users navigate through links, citations, and search results to expand or reorient their exploration (Bates, 1989; Pirolli and Card, 1999, 2005; O’Day and Jeffries, 1993; Huberman et al., 1998). This exploration process helps build contextual understanding, broader topic awareness, and mental maps of the information space (Suh et al., 2023; Bates, 1989; Pirolli and Card, 1999), enabling the users to begin to recognize relationships between concepts, sources, and navigation paths (Aula and Russell, 2008; Teevan et al., 2004). It also creates opportunities for serendipitous discovery, where users encounter unexpected connections and insights (Foster and Ford, 2003; Agarwal, 2015).
Scouting is tightly coupled with filtering (Palani et al., 2022; Pirolli and Card, 1999). Users continuously assess relevance, track what they have seen, and decide what to keep or discard (Marchionini, 2006; Bates, 1989; Chang et al., 2021a). These processes are demanding. Users must scan and manage parallel sessions while distributing attention and memory across scattered information across webpages (Marchionini, 2006; Bates, 1989; Pirolli and Card, 1995; von der Weth and Hauswirth, 2013). Tabbed browsers amplify the burden. They present webpages as a flat, linear stack, requiring users to view them in isolated viewports and switch back and forth to compare and cross-reference (Chang et al., 2021a, b; Dubroy and Balakrishnan, 2010; Brown and Shillner, 1995). Following each page manually can also be costly and discourage exploration (Huberman et al., 1998; Min et al., 2025). As exploration progresses, maintaining context or spotting connections becomes increasingly difficult (Ma et al., 2023; Chang et al., 2021b, a; Weinreich et al., 2006).
Prior work explored visualizing information scent, traces, and relationships among webpages (Cockburn and Greenberg, 2000; Olston and Chi, 2003; Ayers and Stasko, 1995; Lamping et al., 1995; Trigg, 1988; Dörk et al., 2008), helping manage browsing progress (Newfield et al., 1998; Chang et al., 2021b), and providing assistive browsing support, such as recommendations, to promote exploration (Lieberman, 1995; Rhodes and Maes, 2000; Rhodes and Starner, 1996). However, users still have to drill down into each page manually, one by one, to review content and complete navigation, making the process slow and repetitive.
We argue that information and interactions on the web are fundamentally distributed, and the browser must evolve beyond supporting isolated, single-page workflows (Ma et al., 2023; Brown and Shillner, 1995). Instead of leaving users to manually manage and constantly switch among different pages, we explore supporting viewing web content from multiple pages at the same time and operating on them in parallel, to help comparison, maintain context across pages, and follow multiple exploration threads simultaneously.
2.3. Revisiting and Collecting
As exploration grows more complex, the need to revisit information increases, but doing so becomes more demanding and time-consuming. Revisiting information helps refresh memory, preserve context, track updates, and form new connections and ideas (Adar et al., 2008; Tauscher and Greenberg, 1997; Kittur et al., 2013; Capra and Pérez-Quiñones, 2003; Morris et al., 2008; Jones et al., 2001). However, this process can be costly in web browsing. Users may keep related tabs or bookmarks, which quickly accumulate and lead to disorganized spaces and an increasingly difficult refinding process (Chang et al., 2021a, b; Ma et al., 2023). If a webpage is closed, current browsers poorly support refinding visited content, as they only record links and page titles (Chang et al., 2021a).
An important approach to facilitate revisiting is to collect information into a dedicated personal workspace, like a document or folder, and reorganize it according to needs and interpretations (Kittur et al., 2013; Pirolli and Card, 1995, 2005). This process externalizes intermediate thinking and builds structured understandings of the information space (Kittur et al., 2013; Pirolli and Card, 1995, 2005). However, it requires significant manual effort to transform information and restructure the collections, disrupting workflow (Liu et al., 2022b; Kittur et al., 2013).
Prior work has explored supporting the collection of information on the web by enabling the capture of finer-grained units (schraefel et al., 2002; Kittur et al., 2013), reducing collection efforts (Liu et al., 2022b; Stylos et al., 2004; Chang et al., 2016), automating collection or organization (LeeTiernan et al., 2003; Won et al., 2009; Liu et al., 2022a; Chang et al., 2021b), contextualizing exploration with collected information (Palani et al., 2022; Rachatasumrit et al., 2021), and enabling in-situ and progressive sensemaking through growing collections (Liu et al., 2019; Kuznetsov et al., 2022).
In this work, we shift to treating the browser as the primary information workspace, where users can revisit, collect, and organize multiple webpages simultaneously. Webpages can be flexibly structured, extended, and abstracted into simpler formats for faster scanning and easier refining.
2.4. Synthesizing and Transacting
After gathering information, users compare and integrate materials to form a coherent understanding. They organize evidence, generate insights, and identify gaps in their knowledge (Pirolli and Card, 2005). This process helps users refine their mental models and create structured interpretations of the information (Pirolli and Card, 2005; Russell et al., 1993; Pirolli and Card, 1999; Liu et al., 2019).
Synthesizing is cognitively intensive. Users must track multiple representations, manage conflicting or redundant information, and translate unstructured content into coherent frameworks (Pirolli and Card, 2005; Kittur et al., 2013). Synthesizing is often gradual and interleaved with foraging (Pirolli and Card, 1999). However, in traditional web browsing, this task often occurs in a disconnected space from browsing and exploration: users switch to separate tools—documents, spreadsheets, or note-taking apps—to build structured outputs. This fragmentation loses context and interrupts workflows (Kittur et al., 2013; Liu et al., 2022b).
Task-oriented decision-making and operations are increasingly prevalent in web activities (Chang et al., 2021b). Meanwhile, navigating and operating the highly interactive web are essential to many of the previously mentioned activities, from scouting to filtering. However, users often face tedious and repetitive low-level operations that slow down the exploration and hinder flexible synthesis (Bolin et al., 2005; Pu et al., 2023).
Meanwhile, web and UI automation agents powered by LLMs show increasing promise in reducing manual efforts in web browsing (Zhou et al., 2023; Müller and Žunič, 2024; He et al., 2024; Koh et al., 2024; Deng et al., 2023; Huq et al., 2025). These agents, including OpenAI Operator and Claude Computer Use, translate natural language instructions into concrete actions (like clicks and scrolling), continuously analyze unseen interfaces, and adaptively determine the next steps to take without requiring extensive end-user training, scripting, or step-by-step monitoring (ope, 2025; com, 2024). While their performance in the wild still needs substantial improvement, the rapid advancement of LLMs suggests we are stepping into a promising future where automation agents may reliably and swiftly navigate convoluted web environments, gather information, and complete complex tasks.
In this work, we explore on-demand automation and flexible, progressive synthesis. By displaying multiple webpages side by side, users can monitor multiple web automation agents and control repetitive operations across pages in parallel. We also leverage the synthesis capabilities of LLMs to dynamically generate digests that adapt to changing selections and evolving page content, helping users understand information at scale.
3. Summary and Design Goal
Informational tasks on the web involve fluidly interleaved activities from starting and scouting, to synthesizing and transacting. We synthesize key insights from prior literature in Table 1, outlining the typical tasks, user goals, associated costs and benefits, and potential opportunities for facilitation.
Prior research has highlighted the value of manual effort in these activities: It supports anchored exploration aligned with evolving and personal needs (White and Roth, 2009; Ellis, 1989; Teevan et al., 2004); enables serendipitous discovery (Foster and Ford, 2003); and fosters progressive understanding of the problem space (Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005). Manual interaction also contributes to building reusable knowledge, understanding the reasoning behind the final results, and maintaining user agency and trust (Teevan et al., 2004; Pirolli and Card, 1995).
Meanwhile, web-based information tasks require users to repeatedly navigate through distributed webpages and manage large volumes of fragmented information and context. Traditional tabbed browsing, however, prioritizes working within each individual page and barely helps with cross-page activities, amplifying navigation and synthesizing costs (Chang et al., 2021a; Ma et al., 2023; Brown and Shillner, 1995).
Therefore, this work aims to facilitate a user-driven approach to the web information space to preserve the merits of manual effort, while reducing cognitive and interaction costs carried by the distributed nature of web-based tasks.
Figure 2. The Orca interface features a Web Canvas (b), where webpages can be spatially arranged and organized, and a side panel (a), which can be toggled (1) or set to fullscreen (4). The side panel shows pinned pages or summaries—either for individual webpages or multiple selected ones (3). Users can zoom out the viewport to view all pages at once (2). When a webpage is pinned, its instance on the canvas turns purple (6). Users can pin a set of pages and switch among them (5).
3.1. Design Goal: Browsing at Scale
We aim to enable browsing at scale to reduce repetitive low-level operations and constant context switching, supporting tasks across many webpages. Browsing at scale means:
  • $\circ$
    Instead of restarting the search with new queries and sources repeatedly, or navigating into a single thread or subpage at a time, users could start and navigate at scale by expanding from their current context to surface a broad set of relevant resources and exploring multiple pages in parallel.
  • $\circ$
    Instead of clustering pages in linear, cluttered stacks and jumping back and forth between tabs, users could view and organize at scale across multiple pages at the same time.
  • $\circ$
    Instead of finding information from each page individually, users could extract at scale by surfacing key content across many pages simultaneously.
  • $\circ$
    Instead of repeating low-level actions inside individual pages, users could operate at scale by broadcasting intents and operations across multiple pages in one action.
  • $\circ$
    Instead of making sense of information from fragmented sources manually, users could synthesize at scale with summaries that integrate content of multiple sources.
4. Approach
We explore a two-fold design approach to facilitate at-scale, user-driven browsing, as detailed in this section.
4.1. Webpages as Malleable Materials, and Browser as a Malleable Space
We address the limitations of traditional tabbed browsing, which treats each page as an isolated, primary workspace. Tabs are organized in a linear stack, making it difficult to view or operate across multiple pages at once. Additionally, individual pages often contain unrelated content, requiring users to repeatedly sift through noise and perform the same navigation and interaction steps.
Therefore, we reimagine webpages as “malleable materials”—elements that can be viewed at varying scales, transformed, and extracted. Correspondingly, we reconceptualize the browser as a “malleable space”–an environment where these materials can be flexibly composed into custom workspaces. Our approach unfolds across these two dimensions as follows:
Malleable webpage materials: parallel view and customized extraction. To enable at-scale actions across pages, malleable webpages support being viewed at a “zoomed-out” scale, allowing users to monitor and interact with many pages simultaneously without losing context (Section 5.1). To reduce distraction and support focused scanning, content across pages can be selectively extracted—surfacing only the relevant information in a customized view aligned with users’ tasks and informational needs (Section 5.4).
Malleable browser space: dynamic organization and composition. Moving beyond the constraints of linear tabs, malleable webpages can be flexibly reorganized, grouped, and composed into various structures, such as tables. These groupings and structures allow users to distinguish webpages for different tasks, visualize relationships across content, and adapt layouts to fit evolving workflows—forming a malleable, fluid, and task-centric browser environment from previously disconnected webpages (Section 5.2).
4.2. User-Driven and AI-Facilitated Orchestration
On the other hand, we aim to enable browsing at scale through a user-driven approach that preserves cognitive benefits and maintains user control over evolving goals and preferences. We explore interface designs and interaction techniques that empower users to navigate, operate on, and synthesize across multiple malleable webpages simultaneously—a process we define as “orchestration.” In this process, AI plays a facilitating role: it accelerates reach, reduces repetitive manual effort, and adapts content to aid sensemaking, while allowing the user to steer and control the collaborative orchestration across malleable pages. Specifically, we explore three collaborative models between the user and AI:
AI scales up, user scales down. To help users quickly bootstrap and expand their exploration across sources, AI interprets the context and content of the pages being browsed and “scales up” the exploration by identifying relevant directions and alternative entry points, and extending the context of available information. The user then “scales down” by reviewing the available materials, filtering based on relevance or their interest, and actively steering the exploration to align with their evolving goals and needs. This collaborative model accelerates the discovery of a broadened information landscape without removing users from the loop, keeping them in control of the direction and depth of exploration (Section 5.3).
AI surfaces, user deep-dives. To support progressive sensemaking that aligns with users’ evolving inquiries, AI helps distill content across malleable webpages into succinct forms and synthesizes cross-source summaries that surface key points in context. Users remain in control of interpretation by selectively deep-diving into areas of interest, posing follow-up questions, and iteratively bringing in new or alternative content for updated extraction and synthesis. This collaborative model supports efficient scanning while preserving direct engagement with source material, enabling a personalized and evolving sensemaking process grounded in intermediate, contextualized digests (Sections 5.45.5).
AI executes, user monitors. To reduce the cognitive and manual overhead of repetitive interactions, AI carries out low-level UI operations in parallel. Users remain in charge by monitoring, guiding, and intervening when needed. They can evaluate results, adjust direction, and override automation in real time. This collaborative model allows users to delegate and expedite tedious tasks while maintaining oversight and control (Section 5.6).
5. Browsing at Scale with Orca
We present our design exploration to enable browsing at scale with our prototype browser, Orca (Figure 2).
5.1. View at Scale: Spatial Overview Across Multiple Webpages
With traditional tabbed browsing, users struggle to constantly jump between multiple linearly stacked tabs to track different browsing threads. To address this, Orca supports showing multiple webpages in parallel by displaying them as “shapes” on the Web Canvas (Figure 2.b). Users can view pages simultaneously and rearrange them freely by dragging them around—much like using a graphics editor such as Figma (fig, 2016). For example, users might cluster pages into separate groups to keep different tasks apart, then cross-reference information or merge groups to work across tasks easily when needed. Unlike prior browsers that organize abstracted webpages spatially (using favicons or titles), Orca keeps webpages fully interactive to support live, cross-page interactions, going beyond earlier spatial tools used for code or writing (Ayers and Stasko, 1995; Bragdon et al., 2010; Jiang et al., 2023a; Suh et al., 2023; Bederson and Hollan, 1994; Agarawala and Balakrishnan, 2006; Henderson and Card, 1986; Wright et al., 2006).
Common graphics editing interaction techniques, such as range selection, also help users reposition multiple pages at once or apply batch operations across selected pages (Masson et al., 2025). Utilizing the spatial layout, users can drag any link from a page onto the canvas to open it at a specific location, while clicking the link opens it next to the original page for easy side-by-side viewing. Orca displays the actual content of multiple pages at once, not just titles or favicons. This helps users quickly get an overview and easily refind pages by recognizing visual cues from the content itself. Users can also zoom in or out, or pan the canvas, to adjust the viewport and focus on different webpages of interest.
As a result, users compose a dynamic, customized workspace at the browser level using multiple webpages. For example, when finding a hotel for a trip, instead of switching back and forth between tabs, users can include multiple pages in view to scan through the differences, move discarded options aside, combine with flight booking webpages and cross-check matches, and zoom in on selected ones to confirm the details.
5.1.1. Pinning webpages for fixed and concentrated views
On the other hand, for any webpage the user wants to view in a fixed manner (e.g., a document they are working on) or highlight, they may pin it to the side (Figure 3.4). Pinned webpages are displayed at the original size, as in a traditional browser, and can be switched among through the Pinned Webpage Bar (Figure 2.a). Unlike the traditional tab bar—where all opened tabs across all tasks are shown, burdening users to organize and relocate them among a large volume of tabs—users may dynamically create the bars with only selected webpages, while others stay open on the canvas.
The pinned sidebar can be used together with the underlying canvas, allowing the user to leverage both the static view of a “base” page and the dynamic, exploratory nature of the Web Canvas. For example, they may pin a webpage of a notetaking app on the side while continuing to explore information using the Web Canvas on the right. The pinned sidebar can also be hidden or expanded fullscreen for a concentrated view of only the pinned pages.
5.2. Organize at Scale: Flexibly Organizing Webpages with Different Structures
Figure 3. In Orca, webpages can be arranged side by side. Users control individual pages via bottom controls—e.g., update URL (2) or navigate back or forward (3)—like in traditional browsers. They can also act on selected pages using the contextual menu (top), indicated by the selection box (8). The viewport can be adjusted to show one or multiple selected pages (1). Pages can be pinned to the sidebar for static view (4), arranged in a grid (5), batch closed (6), or extracted via the Page Extraction toggle to surface key information (7).
Figure 4. Webpages can be organized into a grid (a) or stack (b) layout. Upon creation, Orca assigns the group a short, customizable name based on page content (1), along with its total page count (2). Users can switch between grid and stack layout (3) or convert the group into a table or visualization (4). Pages can be sorted (5) or selectively chosen (6) based on custom criteria. Users can adjust the number of columns in the grid layout to control its width (8). When no longer needed, groups can be dissolved (7).
Manually managing webpages on a freeform canvas and adjusting the viewport can be demanding. Unlike positioning shapes in Figma, where layout is part of the design task itself, this effort can be perceived as an extra burden (Kosch et al., 2023). Meanwhile, a cluttered canvas may make it harder to discover connections, synthesize information, or cross-reference pages. Therefore, Orca offers various ways to help gather and tidy webpages, as well as compose them into different structures and representations.
Orca offers grid and stack organizations of webpages. Users can arrange pages into a grid for a clean layout for quick scanning (Figure 4.a). They can also sort the grid by content or browsing metadata (including time opened or last interacted) to understand relationships and quickly locate webpages. To archive content with minimal canvas space while keeping tabs open, or for pages not currently in use, users can use stacks, where they may flip through the webpages inside (Figure 4.b). Webpages can be dragged in or out of grids and stacks as needed.
Prior work has explored representing multiple webpages as diagrams or nested lists to show relationships among them (Ayers and Stasko, 1995; Chang et al., 2021b). To further support diverse understanding and tasks, Orca explores representing sets of webpages with tables and interactive graphics (Jiang et al., 2023b). Orca displays content from each page in columns of a table. Users can continue adding more columns to extract more information across all pages in the table view. Webpages can also be visualized as interactive diagrams (e.g., bar charts) using Vega-Lite grammar generated on-demand for visualizing a specific aspect of the web content across pages, e.g., when comparing hotel prices (Satyanarayan et al., 2017).
Figure 5. Users can extract and surface information across pages at scale for quick scanning. Users start by selecting a set of webpages (a), entering a simple extraction query (b), and pressing the Extract button (1). Orca then displays the extracted information in a consistent format across pages. Users can continue customizing the views of pages to surface the information they care about (c). Extraction queries are saved (2), and hovering over a query highlights the corresponding extracted information for all pages. Users can further close or sort pages (3) to continue exploration.
Figure 6. Orca facilitates in-depth exploration by allowing users to batch open multiple links from a webpage. The Batch Open menu appears minimized by default at the top right of selected pages (a). On hover, users can pick from suggested queries or enter a custom one (b). Orca then searches the page and compiles matching links, showing the total count to support query refinement (c). Users can drag the capsule onto the Web Canvas to open all links, automatically arranged in a grid (d).
Figure 7. Orca supports in-breadth exploration through Contextual Expansion, which opens relevant pages based on context and user input. When a webpage is selected, a small handle appears at its bottom right corner (a). Clicking it reveals the Contextual Expansion menu, where users can pick suggested queries or enter a custom one to scout related pages (b). Users can drag the handle to spatially set how many pages to add (c), or double-click to let the AI decide. Orca then opens the relevant pages; if a page is not directly accessible via a link, a web automation agent completes the navigation (d).
Figure 8. Multiple automation agents can operate in parallel on their respective webpages, with color-coding for visual tracking and distinction. Each agent has a control bar under the page it is manipulating, showing its overall goal and options to pause or deactivate it (1). A virtual cursor represents the agent and its actions. Clicking the cursor opens a contextual menu to give instructions or fully take over control. When taken over, the virtual cursor snaps to and aligns with the user’s cursor (2). Moving the cursor outside the page returns control to the agent and resumes automation.
Figure 9. The global command bar with feedforward options (right) and the current webpages in the viewport (left). Options of varied types are ordered from bottom to top by relevance for quick access using the arrow keys. (a) After searching for flights to Paris, Orca uses the context to suggest relevant create actions. (b) Organize actions support canvas-wise content search and quick revisiting of related pages. (c) Operate and query actions allow the user to interact with or synthesize across pages. Hovering over an option highlights corresponding webpages on the Web Canvas (1).
5.3. Navigate at Scale: Batch Exploration in Depth and Breadth Through Direct Manipulation
Users often explore both in depth and breadth during an information task, in a highly interleaved manner. However, they currently have to manually click each link on a page to drill down, go back to the base page to repeat the process (e.g., for a search results page); or manually switch between threads and re-enter queries to expand their exploration (e.g., to find results for alternative search queries).
The tedious process and the effort required to manage context across multiple threads often discourage users from exploring as much as they would like or fully following and using multiple threads. Therefore, Orca streamlines exploration in both directions of depth and breadth through direct manipulation.
To help users quickly drill down and overview multiple links on a page, Orca gathers links into meaningful groups for Batch Open, such as “Discounted hotels” on a hotel listing page (Figure 6). Users can quickly open these links by dragging the capsule button onto the canvas. Users can also define custom extraction queries, such as “Hyatt with rating above 9”. Orca extracts links to detail pages that match the query (i.e. Hyatt hotels rated above 9.0) and shows the total number found. The user can then drag the capsule or press Return to open the links on the side of the source page.
To expand the exploration in breadth, Orca helps users search for and visit relevant pages based on their current context. After selecting one or more pages, users can click a small handle at the bottom right corner to open the Contextual Expansion menu (Figure 7). Orca analyzes the content of selected webpage(s) and suggests relevant queries. For example, if the user selects a hotel booking page in Paris, Orca may suggest queries like “Find flights to Paris” to complete the trip planning or “Find the same hotel on other platforms” to compare the prices and options. Users may also input customized queries. Users can then drag the handle (like resizing a shape) to control how many results to open, and thus the complexity of the broad exploration, or press Return (or double-click) to let AI decide the number of pages. When a target page cannot be accessed directly, Orca deploys a web automation agent to navigate from a starting point and reach the desired page, e.g., starting from a hotel booking homepage, automatically filling in trip details, and reaching the desired detail page (Section 5.6).
5.4. Extract at Scale: View Customization Through Continuous Page Extraction
As a trade-off for the zoomed-out overview, each page on the Web Canvas appears smaller to fit within the limited screen space. This means web content is shown at a scale that web designers typically do not intend for users (even with the mobile version).
While our prototype shows that users can still easily read and interact with web content at around one-third scale (on a 14-inch monitor)—fitting up to 9 webpages—it is critical to surface relevant content from each page at a familiar, constant size, regardless of the zoom level of the Web Canvas. More importantly, each page often contains clutter that is unrelated to the user’s current task. Helping users quickly spot relevant content across multiple pages can greatly improve their exploration efficiency.
To support targeted, customized information access, Orca enables extracting content and interactive elements from webpages (Figure 5). Users can select multiple webpages, toggle Page Extraction, and enter a query to pull out relevant information. For example, when comparing hotels, they might ask, “Does it have a pool?” or simply, “Pool”, and Orca uses the LLM to interpret the intent, find answers, and display them in a scannable format. Meanwhile, the source of extracted information is highlighted on the original webpage for the user to inspect when needed. This helps users scan through and filter options in a more customized manner than using the search and filter tools offered by the original websites (Min et al., 2025).
Users can also surface interactive components to perform simple actions across pages directly from the overview. Instead of extracting raw webpage code, the Orca prototype provides buttons, input boxes, and text areas as standardized, reusable UI elements that represent the underlying interactive widgets. This approach enables consistent and reliable UI extraction and generation (Cao et al., 2025). The components remain bi-directionally synced with the original web UI via two mechanisms: content listeners update the extracted UI when the page changes, and web automation agents transmit user actions (such as clicks and inputs) back to the page (Section 5.6).
Table 2. Types of Feedforward Options Supported by the Global Command Bar
Types Explanation Corresponding Activities
[Uncaptioned image] Create Start at scale. Generate multiple starting points in parallel to bootstrap a browsing session (such as alternative search queries or a set of suggested pages) based on the current pages in the viewport and the user’s recent browsing history (Xia et al., 2023). Starting
[Uncaptioned image] Operate Operate at scale. Perform task-specific actions on specific pages, such as filling in forms. When needed, multiple agents may be deployed across webpages to perform tasks simultaneously, e.g., “Navigate to the comment section for each hotel page.” Scouting, Filtering, Collecting, Transacting
[Uncaptioned image] Organize Organize at scale. Arrange webpages in the viewport by sorting (for webpages within grid structures), selecting, closing, or locating specific webpages—either by teleporting to them or bringing them into the current viewport (Figure 9.b). Filtering, Revisiting, Collecting
[Uncaptioned image] Query Synthesize at scale. Ask questions about the webpages in view, enabling the system to retrieve and synthesize information across them. Responses appear in the side panel, where users can continue asking follow-up questions. Synthesizing
5.5. Synthesize at Scale: Dynamic Summary and Information Query
Having to maximize the view of webpages one by one and read through each can significantly slow down the exploration and discourage users from opening and gaining information from the large number of webpages that Orca helps create and manage.
On the other hand, AI-driven information retrieval pipelines like Retrieval-Augmented Generation (RAG), or systems like Perplexity, typically provide a one-pass summary over the gathered resources in the end, despite supporting follow-up queries. In contrast, Orca supports an iterative, contextual, and user-driven process of collecting and making sense of sources. As users scout and select webpages, they can dynamically surface intermediate summaries that evolve with their exploration: by toggling the Exposé sidebar, users can view individual summaries for each page or an aggregated summary across the selected set (updated and cached on selection or web content changes). The granular and dynamic summaries facilitate interleaved sensemaking and information gathering.
Users can continue refining their information retrieval by asking follow-up questions about the selected webpages (or pages in the viewport if none are selected). If the answer is not directly available from the current content, Orca deploys automation agents to the relevant pages to further navigate and search for broader context in response to the question (Section 5.6).
5.6. Operate at Scale: On-Demand, Contextualized, and Parallel Web Automation
Prior approaches to web automation typically involve agents completing entire tasks from start to finish with minimal user involvement, such as generating a full report for a question. However, users often struggle to stay updated with the latest information and progress, and cannot effectively provide intermediate feedback for emerging needs and preferences (Huq et al., 2025). Orca takes a different approach by offering ad hoc and “atomic” automation. Agents act as on-demand components that support specific parts of the workflow, allowing for more flexible, user-driven interaction.
As mentioned in previous sections, automation in Orca supports tasks such as navigating newly opened pages to reach the target page, exploring pages to gather additional information in response to user questions, and carrying out direct operational requests, triggered globally or for one or a few selected webpages. Automation is visualized through virtual cursors—one for each agent—allowing users to track and control their actions in real time (Figure 8).
Web Canvas encodes rich context shared across webpages, both visually for the user to track and as global context for agent utilization. As a result, agents are automatically contextualized without requiring users to enter details manually. For example, if several webpages and a form (to be completed using information from those pages) are placed side by side, selecting all and calling an agent to fill out the form allows it to pull necessary information directly from the neighboring page.
Prior systems based on tabbed browsers typically use a single agent to complete tasks sequentially, often leading to slow progress and poor traceability due to frequent tab switches and disconnected views. In contrast, Orca leverages the Web Canvas as a workspace where multiple processes can run and be visualized in parallel, like dashboards or spreadsheets (where computations are carried out concurrently in space (Dourish, 2017)). For example, to check the distance to the airport for a set of hotels, all detail pages can be automatically opened and organized in a grid structure, where an agent is then deployed to each page to navigate to the right section simultaneously and extract the relevant data to generate the summary.
Since all automation runs in parallel and is shown on the canvas, users can easily monitor, intervene, or assist as needed—for example, to bypass CAPTCHAs or rescue a stuck agent. When control has shifted from the agent to the user, the virtual cursor follows the user cursor and aligns with it to indicate the change (Figure 8.6).
5.7. Supporting Diverse Scalable Instructions with Feedforward Prompting
While many at-scale controls across webpages can be performed through direct manipulation on the Web Canvas, Orca also provides a global command bar that enables users to orchestrate malleable webpages using natural language prompts directly (Figure 9).
The bar can be toggled with Control/Command + T or a button, and functions like a traditional address bar (navigating to a URL or searching a query) when Return is pressed directly.
We synthesize the user workflow identified in the problem framing (Section 2) along with the corresponding aspects required to enable browse at scale, to derive four types of supported cross-page requests: create, operate, organize, and query (Table 2).
A common problem with prompting LLM-assisted actions is the black-box nature of interpretation (Subramonyam et al., 2024; Zamfirescu-Pereira et al., 2023; Min and Xia, 2025). Users often do not know what the model can do—or how it will respond—until after the action is performed. With an identified set of cross-page actions, Orca supports feedforward prompting by translating open-ended user input into a list of options clearly describing specific system actions to trigger. Users can then select from these options, which the AI will carry out.
During prototyping, we found that asking the LLM to suggest and rank actions across all four types in one pass was slow and often missed relevant actions in some types. To address this, Orca sends four parallel requests, one for each type, allowing the LLM to identify actions specific to each. Orca then aggregates the results by filtering out low-confidence actions, sorts the options, and presents the most likely ones for the user to choose from (Appendix A).
6. Implementation
The Orca prototype is an Electron app built with React, where each webpage is loaded in its own webview (OpenJS Foundation and Electron contributors, [n. d.]). The Web Canvas interface is implemented using tldraw (tld, [n. d.]). An M4 Max MacBook Pro with 36GB of unified memory handles up to around 80 webpages before freezing. All LLM-based tasks, including web automation and content summarization, are powered by Claude 3.7 Sonnet (20250219). For web automation, we developed custom HTML distillation and agentic pipelines, following common practices in LLM-based web automation (Zhou et al., 2023; He et al., 2024; Müller and Žunič, 2024). As a result, we can reuse the distilled web content for Page Extraction (Appendix B) and support flexible user intervention during the automated process.
Orca is open-sourced [^2] https://github.com/peilingjiang/orca to facilitate further exploration of at-scale web browsing, and to serve as a testbed for advancing parallel agentic web automation pipelines.
7. Preliminary Evaluation
Evaluating a new browser is challenging due to established user habits with tabbed browsing. A comprehensive understanding of many designs and features requires a longitudinal study with daily use. However, security concerns and the performance of the current prototype make it unsuitable for production and beyond this research’s scope. To assess the usability and effectiveness of the new design concepts and features, we conducted preliminary lab studies. We aim to answer the following research questions:
  1. RQ1
    To what extent does Orca support the users in exploring, making sense of, and retrieving information on the web?
  2. RQ2
    How effectively can the users navigate the webpages on a spatial canvas? What challenges do they encounter?
  3. RQ3
    How effectively do users understand the system status, express their needs, and collaborate with AI?
  4. RQ4
    How is the user experience of user-driven and AI-facilitated browsing compared to AI-driven search engines?
7.1. Participants
We recruited 8 participants (4 female, 4 male, aged 21–27) via internal channels at a university. All participants regularly use generative AI tools like ChatGPT. They also have experience with canvas-based graphics editing tools (e.g., Figma, Adobe Illustrator)—four participants (P1, P3, P5, P6) use these tools frequently (more than a few times per week), while the others (P2, P4, P7, P8) use them less often (less than a few times per month).
7.2. Procedure
The in-person study lasted about 60 minutes per participant, with each participant receiving a 20 USD Amazon Gift Card. The study included the following phases:
Introduction (5 minutes)
Participants received a brief introduction to the research and an overview of Orca, covering the main interface components.
Semi-structured walkthrough and tutorial (30 minutes)
Participants were guided through all features with example tasks and scenarios. They interacted with the system independently with experimenter guidance, familiarizing themselves with Orca’s capabilities before moving on to freeform tasks. During the walkthrough, participants were encouraged to also apply or combine features with other tasks to fully experiment with and understand Orca.
Freeform tasks (15 minutes)
Participants were asked to use Orca to complete a task of their choice.
Questionnaire and interview (10 minutes)
After the tasks, participants completed a 5-point Likert-scale questionnaire evaluating system dimensions and overall experience. A semi-structured interview gathered in-depth qualitative feedback.
7.3. Findings
Overall, participants found the batch operations and the overview and synthesis at scale helpful in supporting their browsing process while allowing them to stay in control. One key finding was that when the cost of navigating, operating, and managing context across pages is reduced, users not only scout and synthesize information more quickly, but also develop a greater “appetite” for exploration and information foraging. Users tend to open more tabs than before, which broadens their context and strengthens the information grounding that supports deeper understanding and better decision-making. We summarize the detailed findings of the questionnaire, interview, and user behavior analysis below.
7.3.1. Easier navigation and inspection at scale stimulates users’ desire for exploration (RQ1)
All participants reported that Orca helped them explore more during information-oriented tasks (3 strongly agree and 5 agree in the Likert-scale survey). They felt navigation and understanding across multiple pages became easier and faster. As P4 noted, “I feel more empowered browsing than with the classic browser… one click could count as many more clicks before.” Participants are “more willing to browse more” (P4), and with Batch Open and Contextual Expansion, they are “able to open [like this] many tabs all at once… [which was] hard to navigate [with traditional browsers], but here you can visually see what each tab is and organize it nicely” (P3) leading them to explore more pages and options.
For example, when choosing a restaurant, P5 opened multiple Yelp lists for different cuisines side by side, used AI to open multiple options from each, and filtered and sorted the pages with Orca to find the best options for each cuisine. “Some wanted Vietnamese food, one wanted Italian. Before, we had to decide the cuisine first, then search,” as restarting separate searches and comparing results were time-consuming, “but now I’m definitely encouraged to search more since I’m not spending any more time and I can gather all the results and interact with them at the same time.”
Scanning through many pages with large amounts of information was also made easier with Page Extraction, which participants unanimously mentioned as the most helpful feature (6 strongly agree, 2 agree). When comparing coffee shops, P1 opened a grid of pages and added multiple abstraction queries, including rating, Wi-Fi, parking, price, hours, and outdoor seating. Orca continuously extracted the data, allowing P1 to scan quickly without distractions: “These are all the things I think about when I look at coffee shops, and I can see them all here without scrolling up and down for each page.” They were then able to quickly filter out places that did not meet their needs. As a result, P1 noted, “I am no longer afraid of opening lots of pages at first, because I know I can then go through them quickly and find what I want.”
7.3.2. Custom use of the spatial layout helps with diverse activities of the exploration process (RQ2)
To our surprise, all participants found the spatial layout of webpages easy to navigate and manage, regardless of their experience with graphics editing tools (5 strongly agree, 3 agree). They appreciated that the Web Canvas provides the “bird’s-eye view” (P7) and “mind mapping” (P8) of multiple websites, which became “notably more helpful as the number of webpages grows” (P6), and helped them stay aware of their exploration progress and the pages they had opened (P1–3, P7).
Participants frequently referred to their experience with graphics editors. For example, P5 described using Figma not just for design work, but to manage all project-related documents and information, where they can easily compare changes across duplicates side by side, instead of “dealing with a bunch of separate Google Docs.” They appreciated that now with a spatial browser and “persistent webpages,” they could also preserve videos and interactive tech demos alongside other content in one workspace. P1 noted that the dynamic, customizable layout of the canvas helped organize messy and divergent informational tasks. They compared it to graphic design, pointing out that “how humans process information and workflows are never necessarily structured.”
We observed that all participants intuitively organized webpages by task, placing them in different areas and forming distinct clusters. P5 liked how they could “assign meaning or context to different areas” without making concrete commitments to folders and structures. For example, after using Orca to surface the win-loss ratios for all thirty NBA teams from their respective homepages, P3 clustered the pages (i.e. teams) into two groups—those with positive and negative ratios. P3 then used the dynamic summary feature to generate summary stats and descriptions for each cluster.
With the help of Orca’s search function, they can then jump between these clusters. For example, while filling out the questionnaire (a Google Form) in Orca, P4–8 used the global command bar to search for and jump to the webpages they wanted to reference while thinking aloud about their ratings. Participants appreciated how spatial separation helped them maintain clean, focused workspaces without needing to close tabs. As P6 said, “I often move the tabs I don’t want to work on now to a separate window… [but] if I [accidentally] close it, I lose all my tabs.”
7.3.3. Contextualized and atomic AI facilitation adapts to diverse preferences for human-AI collaboration (RQ3)
Participants were convinced that Orca effectively tracks their context to provide corresponding assistance (2 strongly agree, 6 agree). P4 was surprised to see feedforward suggestions like “Restaurants in Paris for two people” appear after just typing “Restaurants” in the global command bar, after searching for flights to Paris. However, they were actually looking for local dining options for that evening. By scrolling the viewport to move the flight search pages out of view, Orca updated the suggestions accordingly, removing Paris from the context.
On the other hand, we observed diverse preferences in how participants used AI facilitation features. For example, P1, P2, and P7 utilized features like Batch Open and Contextual Expansion for speeding up exploration, but preferred to read pages themselves from time to time, pinning pages of interest to the side. In contrast, P6 and P8 relied more on summaries and chatting with the AI to obtain information from pages. These differences likely stem from the varied browsing habits people have developed over time (Chang et al., 2021a; Teevan et al., 2004). By integrating AI support into each activity of web browsing and information foraging, from scouting to synthesis, Orca adapts to users’ diverse needs, allowing them to flexibly control certain aspects while delegating others to the AI.
We also observed varied attitudes toward the effectiveness of monitoring and controlling automation agents (2 strongly agree, 1 agree, 3 neutral, 2 disagree). P2 and P7 appreciated being able to leave agents running “in the background” (P2) or “at the periphery” (P7) while continuing their own tasks. P7 noted, “I feel I’m frozen looking at what [OpenAI] Operator is doing… [while] it feels opportunistic [with Orca] in a way that I can do more things while waiting for the agents to complete the task,” and added that it is easy to “keep track of multiple agents at the same time… as long as it’s less than 10.” In contrast, P8 felt “overwhelmed when more than 3 or 4 [agents] were active.” P4 also found the virtual cursors distracting and preferred not to see them as long as the agents completed their tasks correctly. The diversified responses point to the need for further investigation into user preferences and cognitive limits in monitoring and controlling real-time multi-agent systems.
7.3.4. Direct engagement with and control over information sources fosters more personalized information retrieval and strengthens user trust in the results (RQ4)
Participants reported feeling more directly engaged and in control of information and its sources when using Orca compared to AI-driven tools like ChatGPT and Perplexity (2 strongly agree, 6 agree). P6 noted that as the process is more “transparent”—they could see which links the agents opened and verify the content themselves—they “trust the results more.”
P3 emphasized the value of being able to “do their own research” alongside the AI, rather than “blindly” accepting AI-generated summaries. P1 even mentioned that they might not use the summary feature at all, since they now have access to “the actual websites.” Previously, they were often worried that ChatGPT might “miss certain information” and felt “mistrustful” of the results.
P2 appreciated that Orca allowed them to control the sources during the search. For example, when researching synthesizer options, Orca initially opened a set of relevant sites, including reviews, Reddit threads, and shopping pages. P2 then used Contextual Expansion with only Reddit pages to gather more Reddit posts, and selected all of them to generate a summary. As P2 explained, “I know Reddit is my community, and I just want information from there,” while reviews or shopping sites might contain sponsored content. To compare, P2 opened an AI-driven search engine during the study. Despite being explicitly requested, it still generated a summary based on multiple mixed sources.
P4 highlighted the benefits of source-level control in tasks involving personal information and judgment, “I trust Expedia, I trust Airbnb, but not other less common ones—so even if you open a bunch, I’ll only check the ones I trust.”
7.3.5. Limitations
The main limitation of our study lies in the short time participants had to learn and use Orca. By the end of the session, participants had an average of 37 webpages on the Web Canvas. As P5 noted, linear tabbed browsing has become “muscle memory,” and it can feel overwhelming to adjust to a new browsing modality within a limited time frame. Although we did not capture any complaints about the spatial layout of webpages during the study, this may also be because the tasks were relatively contained and did not span long periods, as real-world tasks often do. Participants may not have accumulated a large enough number of pages or complex workflows to fully experience the challenges—or benefits—of navigating a dense canvas across multiple tasks.
A secure, robust, and highly performant implementation is required to further evaluate Orca’s long-term usability in complex and diverse real-world settings.
8. Discussion and Future Work
Reflecting on our design process, study findings, and limitations of the current prototype, we discuss future directions for advancing user-driven, at-scale browsing and other informational tasks.
8.1. Blessings and Curses of Spatial Information Interfaces
Orca enables a parallel overview across webpages through a spatial canvas, while helping users access content from “miniaturized” pages using Page Extraction and dynamic summaries. The canvas representation supports flexible, evolving workflows, allowing users to accumulate and structure content in multi-dimensional yet unconstrained ways (Bederson and Hollan, 1994; Agarawala and Balakrishnan, 2006). Meanwhile, this “freeform, organized chaos” (P5) can lead to visual clutter. Misaligned items and overlapping elements may create extra overhead for users who want a clean, orderly workspace (Kaptelinin and Czerwinski, 2007). In contrast, while linear stacks are semantically messy, they are visually tidy. Future research could explore several directions to augment spatial information interfaces:
First, improve content organization through automatic alignment. While Orca currently supports grid and stack layouts, other structures—such as content-based categorization or localized gestures like snapping based on semantic relationships—could offer more meaningful or lightweight alignment options.
Second, enhance viewport management to reduce the manual effort of panning and zooming. We experimented with rule-based viewport management, where the view could automatically snap to one or a set of pages. However, it failed to deliver significant improvements, as users’ focus shifted quickly during exploration, and rule-based view snapping often disrupted the fluid transitions needed for dynamic tasks. Future work could explore content- and context-aware viewport snapping to better support fluid workflows.
Finally, an overview across web content, or information, does not have to be a spatial one. Future research could investigate alternative structures, such as a linear grid (similar to a spreadsheet), where users can place webpages or other content in fixed cells, to understand the tradeoffs compared to spatial layouts.
8.2. Malleable User Interfaces for the Web
Today’s browsers primarily serve as “renderers” for content and interfaces predefined by website designers and developers, respecting their intended styles, priorities, and even rate limits for how content is consumed. However, these independently designed webpages often fall short in supporting diverse user needs or working seamlessly together for tasks that span across multiple sites (Min et al., 2025). Moreover, interaction mechanisms built for human users are poorly suited for the growing demands of automation agents (Zhou et al., 2023).
Orca treats webpages as malleable materials that users can flexibly overview, represent with different structures, and transform into succinct formats in the malleable browser space. However, webpages are still primarily displayed as isolated “rectangles” that preserve the styles and formatting of the original websites, which are cluttered with irrelevant content—although Page Extraction offers simple information retrieval and UI component surfacing.
With rapidly advancing LLM capabilities and emerging agentic infrastructures, such as Model Context Protocol (MCP) and NLWeb, we anticipate AI agents will soon be able to navigate, operate, interpret, and extract information from most websites with greater reliability and depth (llm, 2024; Anthropic and contributors, 2025; Microsoft, 2025). To this end, we envision treating webpages not merely as visual interfaces, but as “services” or APIs that provide rich data, content, and functionality.
LLM-powered browsers can then serve as interface “composers” and leverage their generative capabilities to create custom, task-specific interfaces layered over the web-served content. Agentic pipelines serve as a transformation layer that connects the custom UI and data of varied formats and schemas (Anthropic and contributors, 2025; Microsoft, 2025). This would enable the creation of malleable, personalized, and evolving workspaces that can selectively incorporate information aligned with users’ tasks across “data sources” (Cao et al., 2025; Min et al., 2023).
8.3. Parallel and Collaborative Real-Time Multi-Agent Systems
Orca introduces more granular, contextualized, and multi-threaded web automation by allowing multiple agents to operate in parallel across pages to complete tasks.
However, feedback from participants revealed a need for better monitoring and control of real-time multi-agent systems. P6 suggested a global progress bar to track agents in one place, rather than scattered across the Web Canvas. P2 proposed breaking down and visualizing the agents’ detailed steps using an interactive diagram to make it easier for users to track and adjust their behavior (Jiang et al., 2023a; Kazemitabaar et al., 2024).
Therefore, future research can explore new mechanisms to monitor agent progress and enable timely user intervention. This could involve methods such as direct manipulation, demonstrations, or even a “group chat” interface for users to give intermediate feedback to one or more agents simultaneously (Lau, 2009).
9. Conclusion
We introduce Orca, a prototype web browser that reconceptualizes webpages as malleable materials and supports user-driven, AI-facilitated orchestration across them in the malleable browser space. With a spatial canvas interface, Orca provides features to enable browsing at scale across webpages, addressing the distributed nature of web-based informational tasks. Our evaluation shows that Orca reduces the effort required to manage complex browsing workflows while preserving user agency, increasing engagement, and stimulating users’ desire for information foraging. These findings highlight a promising direction for future browser designs that tightly integrate AI capabilities into the personal, user-driven informational tasks.
Appendix A Constructing Feedforward Options Lists from Parallel LLM Responses
To reduce a user prompt to concrete actions the system can perform, Orca sends four parallel prompts, one for each category (create, organize, operate, and query), to the LLM. The model returns candidate actions with confidence scores, and the system aggregates them into the feedforward options list as follows.
Let $C=\{c_{1},c_{2},c_{3},c_{4}\}$ be the set of categories. For every $c_{i}$ the LLM yields
$A_{i}=\{\,(a_{ij},\,s_{ij})\mid s_{ij}\in[0,1]\,\}.$
Filtering
Discard low-confidence items and merge the rest:
$A=\bigcup_{i=1}^{4}\bigl{\{}\,(a_{ij},s_{ij})\in A_{i}\mid s_{ij}\geq 0.2\bigr{\}}.$
Sorting
Sort the remaining in descending order of the scores:
$A_{\text{sorted}}=\bigl{[}(a_{1},s_{1}),\dots,(a_{|A|},s_{|A|})\bigr{]}.$
Selecting the final list
Index the list with $k$. Define the running total of confidence scores of selected actions as
$S_{n}=\sum_{k=1}^{n}s_{k},$
and choose the smallest prefix length
$N=\min\bigl{\{}\,n\in\{1,\dots,|A_{\text{sorted}}|\}\,\big{|}\,S_{n}\geq 3\text{ or }n=7\bigr{\}}.$
The final action set is
$A^{*}=\{\,(a_{k},s_{k})=A_{\text{sorted}}[k]\mid 1\leq k\leq N\}.$
Appendix B HTML Distillation for Web Automation and Page Extraction
Following common practices in HTML distillation, we develop a custom distillation algorithm to convert the often highly complex HTML structures of production websites into a concise and manageable form. This distilled representation allows the text-based Orca agents to reason over and act on web content effectively. We incorporate several notable approaches, briefly described below.
We annotate interactive elements with two interaction types: click (such as buttons) and input (for all kinds of <input /> elements, such as text fields or checkboxes). We also include the aria-label or placeholder text when available and applicable. Each interactive element is annotated using the following syntax:
[\$id\$:interactionType (ariaLabel)] content
For example:
[\$m2l\$:click (Send email)] Send
The id is a stable, three-character unique identifier derived from the element’s xpath, ensuring consistent reference across distillations.
We further annotate active elements—typically a button that has just been clicked or an input box that is currently focused—as:
[ACTIVE] $>$ [\$m2l\$:click (Send email)] Send $<$
Changes in elements since the last distillation are shown in a separate prompt following the full current HTML, using a code diff style, such as:
[-] [\$m2l\$:click (Send email)] Send
[+] [\$6jp\$:click] Cancel Send
This helps the LLM focus on parts of the interface that are likely to be most relevant, such as a newly appeared pop-up modal.
Referring to elements for automation agents
Agents can refer to elements using their stable identifiers when executing automation steps. For example:
{ "action": "click",
  "element": "nuu" }
{ "action": "update-value",
  "element": "617",
  "value": "Hello, World!" }
Referring to elements for Page Extraction
Similarly, when an interactive element is extracted from the original page, represented as a standardized UI element, the underlying “target” element is referred to using the identifier as well. For example:
{ "type": "button",
  "title": "Send",
  "target": "m2l" }
That is, a button titled “Send” will be extracted, and clicking it will propagate the click event to the underlying element with id m2l.